OpenStudy (anonymous):

A coin was flipped 150 times. The results of the experiment are shown in the following table: Heads Tails 90 60 Which of the following best describes the experimental probability of getting heads? It is 10% higher than the theoretical probability. It is 10% lower than the theoretical probability. It is equal to the theoretical probability for this data. The experimental probability cannot be concluded from the data in the table.

OpenStudy (xxmuslimxx):

The experimental probability cannot be concluded from the data in the table.

OpenStudy (kropot72):

The experimental probability of heads is 90/150 = 0.6 If the coin was fair, the theoretical probability of heads is 0.5 Therefore the experimental probability of heads is \[\large \frac{0.6-0.5}{0.5}\times\frac{100}{1}=20\ percent\] higher than the theoretical probability of heads. So none of the answer choices is correct.